Internal energy of a gas is given by U=2+3PV.
The molar specific heat of this gas at constant
volume is

Grade:11

1 Answers

Arun
4 years ago
Dear student
 
gamma = (b+1) /b
 
hence gamma = 3+1/3 = 4/3
 
Hope it helps
 
In case of any difficulty, please feel free to ask.
 
Regards
Arun (askIITians forum expert)